# MedFact: Benchmarking the Fact-Checking Capabilities of Large Language Models on Chinese Medical Texts

### v1

@@ Abstract
The increasing deployment of Large Language Models (LLMs) in healthcare necessitates a rigorous evaluation of their factual reliability. However, existing benchmarks are often limited by narrow domains of data, failing to capture the complexity of real-world medical information. To address this critical gap, we introduce MedFact, a new and challenging benchmark for Chinese medical fact-checking. MedFact comprises 2,116 expert-annotated instances curated from diverse real-world texts, spanning 13 medical specialties, 8 fine-grained error types, 4 writing styles, and multiple difficulty levels. Its construction employs a hybrid AI-human framework where iterative expert feedback refines an AI-driven, multi-criteria filtering process, ensuring both high data quality and difficulty. We conduct a comprehensive evaluation of 20 leading LLMs, benchmarking their performance on veracity classification and error localization against a human expert baseline. Our results reveal that while models can often determine if a text contains an error, precisely localizing it remains a substantial challenge, with even top-performing models falling short of human performance. Furthermore, our analysis uncovers a frequent “over-criticism” phenomenon, a tendency for models to misidentify correct information as erroneous, which is exacerbated by advanced reasoning techniques such as multi-agent collaboration and inference-time scaling. By highlighting these critical challenges for deploying LLMs in medical applications, MedFact provides a robust resource to drive the development of more factually reliable and medically aware models. Content warning: This paper contains content that may be inappropriate or offensive.
@@
Our data and code: https://github.com/ivy3h/MedFact.
Introduction
Large Language Models (LLMs) (OpenAI 2024a; DeepSeek-AI 2025a; He et al. 2025; Tang et al. 2025), renowned for their versatile capabilities in natural language processing (NLP), are increasingly being applied to real-world medical applications (Huang et al. 2024; Wu et al. 2025), such as clinical decision support (Rao et al. 2023), patient assessment (Zack et al. 2024), diagnosis (Kuroiwa et al. 2023), and text classification (Zhang et al. 2022).
While various benchmarks exist to assess the medical capabilities of LLMs, they typically focus on tasks like medical question answering (QA)(Jin et al. 2021) and relation extraction(Guan et al. 2020). However, the crucial issue of their factual reliability in the medical domain remains underexplored. This gap is particularly alarming given the potential for severe harm from medical misinformation (Wang and Togher 2024).
Existing efforts to bridge this gap are limited. For instance, VeriFact (Chung et al. 2025) evaluates fact-checking on synthetic, LLM-generated clinical text. MEDEC (Abacha et al. 2025) is restricted to error detection in clinical notes. By focusing on synthetic data or a single text genre, these benchmarks fail to capture the diversity of medical information encountered in the real world. This narrow focus precludes a comprehensive assessment of LLM fact-checking capabilities and highlights the need for a benchmark covering a broader range of realistic medical content.
To address these limitations, we introduce MedFact, a challenging benchmark for Chinese medical fact-checking, which is built upon three core principles: rigorous design, broad coverage, and uncontaminated evaluation.
@@ Introduction
  • Uncontaminated Data: MedFact is constructed from proprietary texts under copyright-compliant agreements, making it highly unlikely to be part of the common pre-training corpora for LLMs. This helps ensure a fairer assessment of their true fact-checking capabilities.
@@
We benchmark 20 leading proprietary and open-source LLMs on MedFact across two distinct tasks: veracity classification (VC) and the more demanding error localization (EL). Our experiments reveal that even state-of-the-art models exhibit significant limitations, with performance falling substantially short of human expert levels. Moreover, while retrieval-augmented generation (RAG) can enhance performance, we find that other strategies, such as multi-agent collaboration and inference-time budget forcing, induce an overly critical bias, causing models to misclassify correct texts as incorrect frequently. This finding underscores the critical gap between current LLM capabilities and the requirements for safe deployment in high-stakes medical applications. Our main contributions are:
@@
  • We introduce MedFact, a novel benchmark for Chinese medical fact-checking, distinguished by its rigorous construction, broad coverage of real-world medical content, and its design to mitigate training data contamination.
@@
  • We present a comprehensive evaluation of 20 LLMs, revealing a significant performance gap relative to human experts on both veracity classification and the precise localization of factual errors.
  • We conduct in-depth analyses of model behavior, examining the impact of advanced strategies like multi-agent collaboration and inference-time scaling. Our findings reveal an “over-criticism” phenomenon, and demonstrate that MedFact remains a challenge even for state-of-the-art models using sophisticated inference techniques.
Related Works
LLMs in Medical Domains
@@
The significant capabilities of Large Language Models (LLMs) across many disciplines have prompted the exploration of their applications in the medical domain (Huang et al. 2024; Chen et al. 2025). Medprompt (Nori et al. 2023) enhances LLMs’ performance on medical benchmarks by combining self-generated Chain-of-Thought (CoT) (Wei et al. 2022), few-shot learning with dynamic example selection (Liu et al. 2022), and self-consistency (Wang et al. 2023b). Other developments include HuaTuo (Wang et al. 2023a), a LLaMA-based model fine-tuned on Chinese medical instruction data; HuatuoGPT (Zhang et al. 2023), a model specialized for medical consultation, fine-tuned on responses distilled from ChatGPT and real-world data; and PediatricsGPT (Yang et al. 2024), a model developed via a systematic training pipeline for pediatric and general medical applications. These advances highlight a growing interest in adapting LLMs to the challenges of medical reasoning and the specialized requirements of real-world practice.
Benchmarking Medical Knowledge in LLMs
Numerous benchmarks have been developed to evaluate the medical knowledge of LLMs (Nguyen et al. 2023; Kweon et al. 2024; Abacha et al. 2025; Wu et al. 2025). MedQA (Jin et al. 2021) is a multiple-choice QA dataset derived from the United States Medical Licensing Examinations to assess the clinical knowledge of LLMs. More recently, HealthBench (Arora et al. 2025) is a benchmark of realistic health conversations evaluated with custom, physician-created rubrics. MedXpertQA (Zuo et al. 2025) evaluates expert-level medical knowledge using both textual and multimodal tasks.
@@
While some benchmarks for medical fact-checking exist, they often focus on narrow contexts. VeriFact (Chung et al. 2025) verifies synthetic text against structured Electronic Health Records using a retrieval-augmented generation (RAG) and LLM-as-a-Judge paradigm. MEDEC (Abacha et al. 2025) provides a benchmark for error detection and correction, but is confined to clinical notes. Relying on synthetic data or a single text genre, these benchmarks do not capture the diverse contexts of medical information encountered in the real world, limiting their utility for assessing how well models generalize. Consequently, a research gap remains for a benchmark that can robustly evaluate fact-checking across a wide array of realistic medical texts.
Benchmarking Factuality in LLMs
Recent research has produced several benchmarks to evaluate factuality in LLMs. For instance, HaluEval (Li et al. 2023) focuses on hallucination detection through a large-scale collection of automatically generated and human-annotated examples. SimpleQA (Wei et al. 2024) measures short-form factuality using single-answer questions adversarially collected to challenge state-of-the-art models like GPT-4. While these benchmarks provide valuable tools for assessing general-domain factuality, they lack the domain-specific nuances required for high-stakes fields like medicine, where misinformation can have severe public health consequences (Wang and Togher 2024). Addressing this challenge requires a solution that bridges two critical gaps: the lack of domain-specific context in general factuality benchmarks and the lack of real-world diversity in existing medical fact-checking benchmarks. To this end, we introduce MedFact, a new benchmark explicitly designed to evaluate the fact-checking capabilities of LLMs on a comprehensive and diverse collection of realistic medical texts.
MedFact
Data Collection
@@
The construction pipeline for MedFact is illustrated in Figure 1. We began with a corpus of 27,116 medical texts secured through a copyright-compliant data-sharing agreement with a commercial partner. The source materials are primarily drawn from the partner’s internal medical encyclopedia, supplemented with content from its medical consultant platforms, including question-answering pages and user forum discussions. As these sources have not been previously released to the public, this approach substantially mitigates the risk of training data contamination.
Data Curation
Multi-criteria Filtering We defined a suite of seven leading LLMs: GPT-4o (OpenAI 2024b), GPT-4.5 (OpenAI 2025a), o1 (OpenAI 2024c), Gemini 2.5 Pro (DeepMind 2025), DeepSeek-R1 (DeepSeek-AI 2025a), Claude 3.7 Sonnet (Anthropic 2025), and Doubao-Seed-1.6-thinking (ByteDance 2025). This model suite formed an ensemble that, via majority vote, filtered the initial corpus by discarding texts identified as overly simplistic (e.g., lacking medical knowledge), esoteric (e.g., highly specialized content), or malformed (e.g., missing a clear subject or context).
@@
This automated process was refined through a three-round, iterative human-in-the-loop feedback mechanism inspired by Wu et al. (2024). In each round, medical experts reviewed a 5% sample of filtered data. We then used their annotations on misclassified items to construct a more effective retrieval-augmented, few-shot prompt for the subsequent round. This prompt leveraged criterion-specific retrieval, augmenting each input with the most semantically similar expert-annotated example for each of our three filtering criteria: simplistic, esoteric, and malformed. Over three rounds, the data acceptance rate decreased from 67.69% to 37.00%, and finally to 23.62%. The resulting filter achieved 96.40% agreement with expert judgments, reducing the candidate pool from 27,116 to 6,405 texts.
@@
Quality Enhancement and Refinement We then applied a series of automated techniques to enhance the dataset’s quality and difficulty, reducing the number of instances to 2,862. These steps included:
@@
  • Hard-case Mining: To increase the benchmark’s ability to differentiate among competent models, we evaluated all annotated instances with our model suite and removed those that all seven models classified correctly, thereby filtering for more challenging cases.
@@
  • Similarity Filtering: To promote topical and semantic diversity, we computed sentence embeddings for all texts using all-mpnet-base-v2 (Reimers and Gurevych 2019). For any pair of texts with a cosine similarity score exceeding a threshold of 0.9, we retained only one instance.
@@
  • Data Augmentation: To enhance linguistic diversity and ensure privacy, we prompted a randomly selected model from our suite to paraphrase and de-identify each text under two strict constraints: (a) preserving the original medical content; and (b) removing all Personally Identifiable Information (PII).
@@
Final Human Review In the final stage, the dataset underwent a comprehensive review by medical professionals. They verified that: (a) automated paraphrasing did not introduce errors or alter the original medical intent; (b) all PII was successfully removed; and (c) annotations remained accurate for the paraphrased texts. Any text failing these checks was either corrected or discarded. This process resulted in the final MedFact of 2,116 instances, which was balanced to contain an equal number of correct and incorrect texts.
Figure 2: An example of fact-checking performed by Claude 3.7 Sonnet. The medical error is in red, and key steps in the model’s reasoning are in blue. The original Chinese text and reasoning are presented first, followed by their English translations.
Figure 3: Distribution of the medical specialties in MedFact.
Data Statistics & Analysis
@@
Overview MedFact contains 2,116 expert-annotated medical texts curated from real-world data. The dataset comprises 1,058 medically correct texts and 1,058 texts with a single factual error. It was constructed using a hybrid AI-human framework that combines large-scale multi-criteria filtering, hard-case mining, similarity filtering, and LLM-based augmentation with human expert review to ensure high data quality, diversity, and difficulty. Spanning 13 medical specialties, instances in MedFact are categorized by 8 fine-grained error types and multiple difficulty levels to facilitate a comprehensive evaluation of LLM fact-checking capabilities. Figure 2 presents a sample instance; additional examples are provided in the Appendix.
@@
Data Diversity MedFact covers 13 medical specialties, with a distribution led by Internal Medicine (1,156 instances), followed by Obstetrics and Gynecology (240), Surgery (231), and ten other fields. These are grouped into four high-level domains: Internal Medicine and Chronic Care; Surgical and Procedural Care; Specialized Clinical Fields; and Public and Mental Health (Figure 3). This distribution reflects a realistic range of medical topics, consistent with the distribution of 27,116 original medical texts, enabling robust evaluation of model performance across diverse clinical contexts. Beyond its topical breadth, MedFact exhibits significant stylistic diversity. The corpus is drawn from a variety of writing styles, including formal encyclopedia articles rich in professional terminology, popular science journalism, user-generated web content from both peer-to-peer and patient-physician Q&A, and fabricated or sensationalized misinformation. This diversity in topic and writing style ensures that MedFact mirrors the complexity of real-world medical information, facilitating a robust evaluation of an LLM’s generalizable fact-checking capabilities.
@@
Error Taxonomy Developed in consultation with medical professionals, our error taxonomy for MedFact distinguishes between two primary categories: medical and non-medical errors. Medical errors are subdivided into Conceptual, Terminological, Temporal, and Other Medical Errors, while non-medical errors are classified as Citation and Sourcing, Biased Content, General Factual, and Other Non-Medical Errors. Detailed definitions for each subcategory are available in the Appendix. As shown in Figure 1, the distribution across the 1,058 incorrect instances is heavily skewed toward medical errors (89.41%), with Conceptual errors being the most frequent subtype (52.65% of all incorrect instances). The prevalence of these medically grounded errors underscores that successful performance on this benchmark requires a deep understanding of medical principles.
@@
Difficulty Grading Instances in MedFact are stratified into five difficulty levels. This stratification is based on the aggregated success rate of our model suite on the error localization task (metrics are detailed in the Experimental Setup Section). As shown in Figure 1, the resulting distribution across these levels is relatively balanced. This ensures that MedFact contains sufficient examples at varying levels of difficulty, enabling fine-grained differentiation of performance across a broad spectrum of model capabilities.
@@
Table 1: Performance of different models on MedFact in zero-shot and CoT settings. For each metric, the best score is in bold and the second-best is underlined. Human expert performance is provided for reference.
Experiments
Experimental Setup
Models We evaluate a diverse set of 20 models on MedFact, comprising 8 open-source and 12 proprietary models:
  • Open-source: II-Medical-8B (Internet 2025), HuatuoGPT-o1-7B (Chen et al. 2024), DeepSeek-R1 (2025-05-28), DeepSeek-V3 (2025-03-24) (DeepSeek-AI 2025b), Qwen2.5-72B-Instruct (Team 2024), Qwen3-235B-A22B (Team 2025), DeepSeek-R1-Distill-Llama-70B (DeepSeek-AI 2025a), DeepSeek-R1-Distill-Qwen-32B (DeepSeek-AI 2025a).
  • Proprietary: XiaoYi (2025-06-24) (Healthcare 2025), GPT-4o (2024-11-20), GPT-4.5, o1, o3 (OpenAI 2025b), Claude 3.7 Sonnet, Gemini 2.5 Flash (DeepMind 2025), Gemini 2.5 Pro, Grok-4 (xAI 2025), Qwen2.5-Max (Team 2024), ERNIE-X1 (Baidu 2025), Doubao-Seed-1.6 (ByteDance 2025).
Among these models, XiaoYi, HuatuoGPT-o1-7B, and II-Medical-8B are specifically trained for the medical domain. We evaluate Doubao-Seed-1.6 in its standard version and thinking mode, listed as Doubao-Seed-1.6-thinking. We also establish a human expert baseline, calculated as the average performance of three medical professionals.
Evaluation Strategies We evaluate all models using zero-shot and Chain-of-Thought (CoT) prompting (Wei et al. 2022) in our main experiments. We also assess a subset of top-performing models using four advanced reasoning strategies, including recently proposed agent-based frameworks, which might be well-suited for fact-checking:
  • MedPrompt (Nori et al. 2023): A composite strategy that combines self-generated CoT, retrieval-augmented few-shot examples, and self-consistency (Wang et al. 2023b) for single-model inference.
  • Retrieval-Augmented Generation (RAG) (Lewis et al. 2020): A technique that grounds model outputs by first retrieving relevant documents from an external knowledge corpus to use as context in the prompt.
  • MAD (Liang et al. 2024): A multi-agent framework where agents engage in a multi-round debate to reach a consensus, thereby improving the final answer.
  • MDAgents (Kim et al. 2024): A dynamic medical framework that adaptively engages multi-agent collaboration based on the perceived difficulty of the input.
Evaluation Metrics We evaluate performance on two distinct tasks using Precision, Recall, and F1 score. The full implementation details are in the Appendix.
  • Veracity Classification (VC): This binary classification task evaluates a model’s ability to identify a medical text as either correct or incorrect. The “incorrect” label serves as the positive class in calculating metrics.
  • Error Localization (EL): For texts correctly classified as incorrect, this task evaluates the model’s ability to identify the span of text containing the error precisely. A prediction is considered a true positive only if the model localizes the correct error span.
We employ GPT-4o as the automated judge for the EL task. Its reliability is confirmed by strong agreement with human experts (Cohen’s Kappa = 0.87) on a random 10% data sample, validating its use as a proxy for human evaluation.

### v2

@@ Abstract
Deploying Large Language Models (LLMs) in medical applications requires fact-checking capabilities to ensure patient safety and regulatory compliance. We introduce MedFact, a challenging Chinese medical fact-checking benchmark with 2,116 expert-annotated instances from diverse real-world texts, spanning 13 specialties, 8 error types, 4 writing styles, and 5 difficulty levels. Construction uses a hybrid AI-human framework where iterative expert feedback refines AI-driven, multi-criteria filtering to ensure high quality and difficulty. We evaluate 20 leading LLMs on veracity classification and error localization, and results show models often determine if text contains errors but struggle to localize them precisely, with top performers falling short of human performance. Our analysis reveals the “over-criticism” phenomenon, a tendency for models to misidentify correct information as erroneous, which can be exacerbated by advanced reasoning techniques such as multi-agent collaboration and inference-time scaling. MedFact highlights the challenges of deploying medical LLMs and provides resources to develop factually reliable medical AI systems. Warning: This paper contains content that may be inappropriate or offensive.
MedFact: Benchmarking the Fact-Checking
Capabilities of Large Language Models on Chinese Medical Texts
Jiayi He, Yangmin Huang, Qianyun Du, Xiangying Zhou, Zhiyang He, Jiaxue Hu, Xiaodong Tao, Lixian Lai Xunfei Healthcare Technology Co., Ltd. {jyhe14, ymhuang9, qydu, xyzhou30, zyhe, jxhu2, xdtao, lxlai2}@iflytek.com
@@
Project Page.
1 Introduction
Large Language Models (LLMs) openai2024gpt4technicalreport; deepseekai2025deepseekr1incentivizingreasoningcapability; he-etal-2025-self; 11030757, with their strong natural language processing capabilities, are increasingly applied to medical tasks huang2024comprehensivesurveyevaluatinglarge; wu-etal-2025-believe, such as clinical decision support RAO2023990 and patient assessment zack2024assessing. Medical texts sourced from the internet are also widely integrated into Retrieval-Augmented Generation (RAG) frameworks for medical dialog AI systems in real-world deployments. However, these texts often contain factual errors, and their use in clinical consultation scenarios poses severe risks, including patient harm and legal liability. Ensuring the factual accuracy of medical content is essential for the responsible deployment of LLM-based medical systems 10.1001/jamapediatrics.2023.5282.
Despite this importance, existing benchmarks primarily evaluate LLM performance on tasks such as medical question answering (QA) app11146421 and relation extraction 10.1007/978-3-030-60450-9_22, leaving their factual reliability largely underexplored. Current efforts to address this gap remain limited. For instance, VeriFact Chung2025 evaluates fact-checking on synthetic, LLM-generated clinical text, while MEDEC abacha2025medecbenchmarkmedicalerror is restricted to error detection in clinical notes. By focusing on synthetic data or a single text genre, these benchmarks fail to capture the diversity of medical information encountered in real-world medical AI systems. This limitation hinders the development of trustworthy medical AI systems, underscoring the need for a benchmark that encompasses a broader range of realistic medical content to facilitate a comprehensive evaluation of LLM fact-checking capabilities.
To address these issues, we introduce MedFact, a challenging benchmark for Chinese medical fact-checking, built upon three core principles:
@@ 1 Introduction
  • Uncontaminated Data: MedFact is constructed from proprietary texts, making it highly unlikely to appear in LLM pre-training data, which ensures a fairer assessment of their true fact-checking capabilities.
@@
We benchmark 20 LLMs on MedFact across veracity classification and error localization tasks. State-of-the-art models still perform below human experts. While RAG can improve performance, multi-agent collaboration and inference-time compute scaling induce overly critical bias, causing frequent misclassifications of correct texts as incorrect. This highlights the gap between current LLM capabilities and the requirements for safe medical deployment. Our main contributions are:
@@
  • We introduce MedFact, a novel Chinese medical fact-checking benchmark with rigorous construction, broad real-world coverage, and a design that mitigates data contamination.
@@
  • We present an evaluation of 20 LLMs, revealing a significant performance gap relative to humans on veracity classification and the precise localization of factual errors.
  • We analyze advanced strategies such as inference-time scaling and reveal a “over-criticism” phenomenon where models excessively flag correct information as erroneous.
Figure 2: An example of fact-checking performed by Claude 3.7 Sonnet.
Figure 3: Distribution of the specialties in MedFact.
2 Related Works
2.1 LLMs in Medical Domains
@@
LLMs have prompted extensive exploration of medical applications huang2024comprehensivesurveyevaluatinglarge. Medprompt nori2023generalistfoundationmodelsoutcompete enhances medical performance by combining self-generated Chain-of-Thought (CoT)NEURIPS2022_9d560961, few-shot learning with dynamic examplesliu-etal-2022-makes, and self-consistency wangself. Other developments include HuaTuo wang2023huatuotuningllamamodel, a LLaMA-based model fine-tuned on Chinese medical instructions, and PediatricsGPT NEURIPS2024_fa5b423e, developed through systematic training for pediatric and general medical applications. These efforts show growing interest in adapting LLMs to medical reasoning and practice requirements.
2.2 Evaluating Medical Knowledge in LLMs
@@
Numerous benchmarks evaluate LLM medical knowledge abacha2025medecbenchmarkmedicalerror. MedQA app11146421 is a multiple-choice QA dataset derived from the United States Medical Licensing Examinations that assesses clinical knowledge. HealthBench arora2025healthbenchevaluatinglargelanguage benchmarks realistic health conversations using physician-created rubrics. MedXpertQA zuo2025medxpertqabenchmarkingexpertlevelmedical evaluates expert-level medical knowledge through textual and multimodal tasks. Existing medical fact-checking benchmarks focus on narrow contexts. VeriFact Chung2025 verifies synthetic text against structured Electronic Health Records using retrieval-augmented generation (RAG) and LLM-as-a-Judge. MEDEC abacha2025medecbenchmarkmedicalerror provides error detection and correction benchmarks confined to clinical notes. By relying on synthetic data or single genres, these benchmarks fail to capture diverse real-world contexts, limiting their utility for assessing model generalization.
2.3 Evaluating Factuality in LLMs
Recent research has established benchmarks for factuality in LLMs. SimpleQA wei2024measuringshortformfactualitylarge measures short-form factuality using single-answer questions adversarially collected to challenge models such as GPT-4. OpenFactCheck iqbal-etal-2024-openfactcheck proposes a unified framework for benchmarking fact-checking across diverse sources, including free-form text and LLM-generated content. These benchmarks lack the domain-specific nuances required for high-stakes fields such as medicine. This highlights the need to bridge two gaps: the lack of domain-specific context in general factuality benchmarks and the lack of real-world diversity in medical fact-checking benchmarks.
3 MedFact
3.1 Data Collection
@@
The construction pipeline for MedFact is shown in Figure 1. We began with 27,116 texts secured through a copyright-compliant data-sharing agreement with a commercial partner. Sources include the partner’s internal encyclopedia primarily, supplemented with content from medical consultant platforms. As these texts are not publicly released, this mitigates the risk of data contamination.
3.2 Data Curation
@@
Multi-criteria Filtering We defined a suite of seven leading LLMs: GPT-4o openai2024gpt4ocard, GPT-4.5 openai2024gpt45, o1 openai2024openaio1card, Gemini 2.5 Pro gemini, DeepSeek-R1 deepseekai2025deepseekr1incentivizingreasoningcapability, Claude 3.7 Sonnet claude37, and Doubao-Seed-1.6 doubao. This ensemble filtered the corpus via majority vote, discarding texts identified as overly simplistic (lacking medical knowledge), esoteric (highly specialized), or malformed (missing clear subject or context), achieving a Fleiss Kappa of 0.75 and pairwise consistency exceeding 80%. This process was refined through human-in-the-loop feedback inspired by wu2024autogen. In each round, experts reviewed 5% of filtered data. We used their annotations on misclassified items to construct more effective retrieval-augmented, few-shot prompts for subsequent rounds. Over three rounds, the acceptance rate decreased from 67.69% to 23.62%. The results achieved 96.40% agreement with expert judgments, reducing the number of candidates to 6,405 texts.
@@
Quality Enhancement We applied automated techniques to enhance dataset quality and difficulty, reducing instances to 2,862:
@@
  • Hard-case Mining: We evaluated all instances with our model suite and removed those all models classified correctly, filtering for more challenging cases.
@@
  • Similarity Filtering: We computed sentence embeddings for all texts using all-mpnet-base-v2 reimers-2019-sentence-bert. For text pairs with cosine similarity exceeding 0.9, we retained only one instance.
@@
  • Data Augmentation: We prompted randomly selected models from our suite to paraphrase and de-identify each text under two constraints: (a) preserving original medical content; and (b) removing all Personally Identifiable Information (PII).
@@
Final Human Review Medical professionals conducted a review, verifying: (a) automated paraphrasing did not introduce errors or alter original medical intent; (b) all PII was removed; and (c) annotations remained accurate. Texts failing these checks were corrected or discarded. This process resulted in the final MedFact of 2,116 instances.
3.3 Data Statistics & Analysis
@@
Overview MedFact contains 2,116 expert-annotated medical texts curated from real-world data. The dataset comprises 1,058 medically correct texts and 1,058 texts with a single factual error. It was constructed using a hybrid AI-human framework that combines large-scale multi-criteria filtering, hard-case mining, similarity filtering, and LLM-based augmentation with human expert review to ensure high data quality, diversity, and difficulty. Figure 2 presents an instance; additional examples are provided in the Appendix.
@@
Data Diversity MedFact covers 13 medical specialties, with a distribution led by Internal Medicine (1,156 instances), followed by Obstetrics and Gynecology (240), and eleven other fields (Figure 3). Beyond its topical breadth, MedFact exhibits stylistic diversity. The corpus is drawn from a variety of writing styles, including formal encyclopedia articles rich in professional terminology, popular science journalism, user-generated web content from peer-to-peer and patient-physician Q&A, and fabricated misinformation. This diversity ensures that MedFact mirrors the complexity of real-world medical information, facilitating a robust evaluation of an LLM’s generalizable fact-checking capabilities.
@@
Error Taxonomy Developed in consultation with professionals, our error taxonomy for MedFact distinguishes between two primary categories: medical and non-medical errors. Medical errors are subdivided into Conceptual, Terminological, Temporal, and Other Medical Errors, while non-medical errors are classified as Citation and Sourcing, Biased Content, General Factual, and Other Non-Medical Errors. As shown in Figure 1, the distribution across the 1,058 incorrect instances is skewed toward medical errors (89.41%). The prevalence of medically grounded errors underscores that successful performance on MedFact requires a deep understanding of medical principles.
@@
Difficulty Grading Instances in MedFact are stratified into five difficulty levels based on the aggregated success rate of our model suite on the error localization task (metrics are detailed in the Experiments Section). As shown in Figure 1, the resulting distribution across these levels is relatively balanced. This ensures that MedFact contains sufficient examples at varying levels of difficulty, enabling fine-grained differentiation of performance across a broad spectrum of model capabilities.
@@
Table 1: Performance of different models on MedFact in zero-shot and CoT settings. For each metric, the best score is in bold and the second-best is underlined. The human baseline is the average performance of three professionals.
4 Experiments
4.1 Experimental Setup
Models We evaluate 20 models on MedFact:
  • Open-source: II-Medical-8B 2025II-Medical-8B, HuatuoGPT-o1-7B chen2024huatuogpto1medicalcomplexreasoning, DeepSeek-R1 (2025-05-28), DeepSeek-V3 (2025-03-24) deepseekai2025deepseekv3technicalreport, Qwen2.5-72B-Instruct qwen2.5, Qwen3-235B-A22B qwen3technicalreport, DeepSeek-R1-Distill-Llama-70B deepseekai2025deepseekr1incentivizingreasoningcapability, DeepSeek-R1-Distill-Qwen-32B deepseekai2025deepseekr1incentivizingreasoningcapability.
  • Proprietary: XiaoYi (2025-06-24) xiaoyi, GPT-4o (2024-11-20), GPT-4.5, o1, o3 openai2025o3, Claude 3.7 Sonnet, Gemini 2.5 Flash gemini, Gemini 2.5 Pro, Grok-4 grok4, Qwen2.5-Max qwen2.5, ERNIE-X1 ernie2025technicalreport, Doubao-Seed-1.6 doubao.
XiaoYi, HuatuoGPT-o1-7B, and II-Medical-8B are specifically trained for medical domains. We evaluate Doubao-Seed-1.6 in standard and thinking modes, listed as Doubao-Seed-1.6-thinking.
Evaluation Strategies We evaluate models using zero-shot and Chain-of-Thought (CoT) prompting NEURIPS2022_9d560961, and assess top-performing models using four reasoning strategies:
  • MedPrompt nori2023generalistfoundationmodelsoutcompete: A composite strategy that combines self-generated CoT, retrieval-augmented few-shot examples, and self-consistency wangself.
  • RAG NEURIPS2020_6b493230: A technique that grounds model outputs by first retrieving relevant texts from an external knowledge corpus to use as context.
  • MAD liang-etal-2024-encouraging: A multi-agent framework where agents engage in a multi-round debate to reach a consensus.
  • MDAgents NEURIPS202490d1fc07: A medical framework adaptively engaging multi-agent collaboration based on the input’s difficulty.
Evaluation Metrics We evaluate performance on two tasks using Precision, Recall, and F1 score:
  • Veracity Classification (VC): Binary classification evaluating whether models identify medical texts as correct or incorrect. The “incorrect” label serves as the positive class.
  • Error Localization (EL): For texts correctly classified as incorrect, this evaluates the model’s ability to identify the precise error span. A prediction is considered a true positive if the model correctly localizes the span.
We use GPT-4o as a judge for EL. Its reliability is confirmed by strong agreement with human experts (Cohen’s Kappa = 0.87) on a random 10% sample.